a cube made of an unknown material has a dimension of 6cm on all edges . if the upper face is displaced laterally by 0.1 mm by a force of 2 x 10 ^6 dynes acting parallel to the surface , calculate the shear modulus of the material


1
Expert's answer
2022-03-11T08:28:00-0500

Answer

shear modulus of the material

"\gamma=\frac{shear\space stress}{shear strain}\\=\frac{F/A}{l/L}\\=\frac{20/0.0036}{1/600}\\=3.33MPa"